OpenStudy (anonymous):

HOW TO FIND THE RATIO OF AREAS:49:16 I need to write the ratio of the lengths of corresponding sides>

OpenStudy (unklerhaukus):

length is square root of the area

OpenStudy (anonymous):

Square root both of them?

OpenStudy (anonymous):

Please help

OpenStudy (unklerhaukus):

7:4

OpenStudy (unklerhaukus):

you had that already right?

OpenStudy (anonymous):

Can you answer this too? doubling the side length of a square always? doubles the area

OpenStudy (unklerhaukus):

well if the side length = s the area = s^2 doubling the side length d=2s the area becomes d^2=(2s)^2=4s^2 which is more than double s^2

OpenStudy (saifoo.khan):

formula is, \[\frac{A_1}{A_2}= (\frac{L_1}{L_2})^2\]